# Inversion sequences growing on the left

### v1

@@ 4 Inversion sequences avoiding 010 and 102
Recall that the generating tree growing on the left construction as we described it in Section 2 cannot be applied to the class $\mathcal{I}(010,102)$, due to the avoidance of the pattern 010. More precisely, applying this construction to every sequence in $\mathcal{I}_{n}(010,102)$ generates every sequence of $\mathcal{I}_{n+1}(010,102)$ in which the value 1 does not appear to the right of a greater value. In order to generate every sequence of $\mathcal{I}(010,102)$, we need to complete this construction with new steps which enable the value 1 to appear at any position (while avoiding the patterns 010 and 102).
The most obvious solution is to simply insert new entries of value 1. Usually, inserting a new entry before the end of an inversion sequence results in a construction which is not surjective. Indeed, if a sequence $\tau$ is obtained by inserting a new entry of value 1 at position $i$ in an inversion sequence $\sigma\in\mathcal{I}_{n}$, then for all $j\in[i+1,n+1]$, we have $\tau_{j}=\sigma_{j-1}<j-1$. In particular, a sequence $\tau$ which contains a saturated entry $\tau_{j}=j-1$ for some $j>i$ cannot be generated this way. However in our case, we are interested in generating a sequence $\tau$ which avoids 102, and in which the value 1 at position $i$ appears to the right of a value $k>1$. In particular, for all $j>i$, $\tau$ must satisfy $\tau_{j}\leqslant k<i-1<j-1$, so the surjectivity problem disappears.
It can be shown that the construction described above indeed defines a generating tree for $\mathcal{I}(010,102)$. In this tree, the children of a sequence are obtained by either applying the same steps as for the generating tree growing on the left, or inserting a new entry of value 1 (to the right of a greater value). In other words, the parent of a nonempty sequence is obtained by
@@
  • removing the first zero and subtracting 1 from all nonzero entries, otherwise.
@@
We introduce a second small change in the construction, to make it more convenient later. If a nonempty sequence contains an entry of value 1 with no greater value to its left (i.e. the sequence begins with a prefix of zeros, followed by a 1), then its parent is now obtained by subtracting 1 from the leftmost entry of value 1 only [^3] When exploring the ascendants of a sequence of this form, its parent in our original definition will eventually appear, after all entries of value 1 have been turned into zeros. We have simply placed some intermediate sequences on that path.. Otherwise, its parent is determined as described just above. The corresponding tree is represented in Figure 2. It should be noted that this tree does not quite conform to the usual definition of a generating tree, as some nodes are on the same level as their parent.
@@
The article [JumpingGT] introduces the notion of ”jumping” generating trees, in which each edge has a (bounded) positive integer length. If a node at level $n$ has a child through an edge of length $k$, then this child is at level $n+k$ (in a classical generating tree, all edges have length $1$). The tree in Figure 2 could be seen as a jumping generating tree having edges of length 0 or 1.
This generating tree allows us to compute the generating function of $\mathcal{I}(010,102)$. However, this result is easier to obtain by restricting our study to the subset of sequences of $\mathcal{I}(010,102)$ which do not contain the value 1. Due to this restriction, the construction presented in the remainder of this section jumps arbitrarily large lengths, so we shall no longer speak of it as a generating tree.
@@
For all $\sigma\in\mathcal{I}$, let $\mathbf{push}(\sigma)=0\cdot(\sigma_{i}+\chi(\sigma_{i}>0))_{i\in[0,n]}$ be the sequence obtained by inserting a zero at the beginning of $\sigma$, and adding 1 to the value of all nonzero entries.
Proposition 6.
The image $\mathbf{push}(\mathcal{I}(010,102))$ is the set $\mathfrak{A}\backslash\{\varepsilon\}$.
@@
Let $\tau\in\mathfrak{A}\backslash\{\epsilon\}$, and let $\sigma$ be the sequence obtained by removing the first zero of $\tau$ and subtracting 1 from the value of all nonzero entries. Then $\sigma\in\mathcal{I}(010,102)$, and $\mathbf{push}(\sigma)=\tau$. ∎
@@
Let $n\geqslant 0$. The function $\mathbf{push}$ is injective, so it describes a bijection between $\mathcal{I}_{n}(010,102)$ and $\mathbf{push}(\mathcal{I}_{n}(010,102))$. By construction, the image under $\mathbf{push}$ of a sequence of length $n$ is a sequence of length $n+1$, so it follows from Proposition 6 that $\mathbf{push}(\mathcal{I}_{n}(010,102))=\mathfrak{A}_{n+1}$. ∎
@@
If $\sigma$ already contains the value 1, then a new entry of value 1 can only be inserted next to the current entries of value 1, by Proposition 5, and we have little interest in the active sites of $\sigma$ in that case. We next characterize the active sites of $\sigma$ when $\sigma$ does not contain the value 1.
@@
The value 1 cannot be inserted at positions $[2,z]$, since it would create a subsequence $(0,1,0)$. Inserting the value 1 at any other position (i.e. in $[z+1,n+1])$ cannot create an occurrence of 010, since there are no zeros to the right and no other entry of value $1$ in $\sigma$. It follows that $\mathbf{Sites}(\sigma)$ is the subset of $[z+1,n+1]$ of positions where inserting the value 1 does not create an occurrence of 102. Note that all values of $(\sigma_{i})_{i\in[z+1,n]}$ are greater than 1, so the new entry of value 1 could only take the role of the ’0’ of the pattern 102. In other words, $i\in[z+1,n+1]$ is an active site of $\sigma$ if and only if no value of $(\sigma_{j})_{j\in[z+1,i-1]}$ is less than a value of $(\sigma_{j})_{j\in[i,n]}$. ∎
Let $\mathfrak{B}=\mathfrak{A}\backslash\{(0)^{n}\;:\;n\geqslant 0\}$ be the subset of non-constant sequences in $\mathfrak{A}$. Equivalently, $\mathfrak{B}$ is the set of inversion sequences which avoid 010 and 102, do not contain the value 1, and contain a value greater than 1. For all $n,z,s\geqslant 0$, let $\mathfrak{B}_{n}=\{\sigma\in\mathfrak{B}\;:\;|\sigma|=n\}$, let $\mathfrak{B}_{n,z}=\{\sigma\in\mathfrak{B}_{n}\;:\;|\mathbf{Zeros}(\sigma)|=z\}$, and let $\mathfrak{B}_{n,z,s}=\{\sigma\in\mathfrak{B}_{n,z}\;:\;|\mathbf{Sites}(\sigma)|=s\}$. Let also $\mathfrak{b}_{n}=|\mathfrak{B}_{n}|$, $\mathfrak{b}_{n,z}=|\mathfrak{B}_{n,z}|$, and $\mathfrak{b}_{n,z,s}=|\mathfrak{B}_{n,z,s}|$.
@@
Let $\sigma^{\prime}=\mathbf{push}^{-1}(\sigma)\in\mathcal{I}_{n-1}(010,102)$ be the sequence obtained by removing the first zero from $\sigma$, and subtracting 1 from every nonzero entry.
@@
    • If $\sigma^{\prime}_{z}=1$ (i.e. the factor of ones appears just after the factor of zeros),
@@
        If $\sigma^{\prime}_{n-1}=1$, then $\sigma^{\prime}=(0)^{z-1}\cdot(1)^{n-z}$ by Proposition 5, so $\sigma=(0)^{z}\cdot(2)^{n-z}$. In that case, $\mathbf{Sites}(\sigma)=[z+1,n+1]$, so $z+s=n+1$.
@@
        Otherwise, $\sigma^{\prime}_{n-1}>1$. In that case, we have $\sigma_{z+1}=2$ and $\sigma_{n}>2$, hence inserting the value 1 at any position in $[z+2,n]$ would create an occurrence of 102. This means that $\mathbf{Sites}(\sigma)=\{z+1,n+1\}$, so $s=2$.
@@
        Since the factor of ones of $\sigma^{\prime}$ is consecutive to its factor of zeros, replacing all the ones by zeros yields a sequence $\tau$ which still avoids 010 and 102. Notice that $\tau$ does not contain the value 1, but contains a value greater than 1, so $\tau\in\mathfrak{B}_{n-1}$. By construction, the number $z^{\prime}$ of zeros of $\tau$ is at least $z$.
    • If $\sigma^{\prime}_{z}>1$ and $\sigma^{\prime}_{n-1}=1$, let $k$ be the number of entries of value $1$ in $\sigma^{\prime}$, and let $\tau=(\sigma^{\prime}_{i})_{i\in[1,n-k-1]}$ be the sequence obtained by removing all entries of value 1 from $\sigma^{\prime}$. In that case, the rightmost $k$ positions $[n-k+2,n+1]$ are active sites of $\sigma$ which are removed in $\tau$, so $\tau\in\mathfrak{B}_{n-k-1,z-1,s-k}$.
    • If $\sigma^{\prime}_{z}>1$ and $\sigma^{\prime}_{n-1}>1$, let $p\in[z+2,n-1]$ be the position of the leftmost entry of value 2 in $\sigma$ (such an entry exists because $\sigma^{\prime}$ contains the value 1). Then $\{z+1,p,n+1\}\subseteq\mathbf{Sites}(\sigma)$, so $s\geqslant 3$.
      Let $\tau$ be the sequence obtained by removing all entries of value 1 from $\sigma^{\prime}$. Let $n^{\prime}=|\tau|$, and $s^{\prime}=|\mathbf{Sites}(\tau)|$, so that $\tau\in\mathfrak{B}_{n^{\prime},z-1,s^{\prime}}$. By construction, $n^{\prime}\leqslant n-2$. It can be seen that $\{i-1\;:\;i\in\mathbf{Sites}(\sigma)\backslash\{n+1\}\}\sqcup\{n^{\prime}+1\}\subseteq\mathbf{Sites}(\tau)$, so $s^{\prime}\geqslant s$.
@@
To complete the proof, it remains to show that each of the five cases above induces a bijection between the set of sequences $\sigma$ and the set of sequences $\sigma^{\prime}$ or $\tau$. We describe the inverse bijection only for the final case, since the previous cases are simpler.
@@
Let $n\geqslant 3$, $z\geqslant 2$, $s\geqslant 3$, $n^{\prime}\leqslant n-2$, $s^{\prime}\geqslant s$, and let $\tau\in\mathfrak{B}_{n^{\prime},z-1,s^{\prime}}$. Let also $p_{1}<p_{2}<\dots<p_{s^{\prime}}$ be the active sites of $\tau$. Let $\sigma^{\prime}$ be the sequence obtained by inserting $n-n^{\prime}-1$ entries of value 1 at position $p_{s-1}$, and let $\sigma=\mathbf{push}(\sigma^{\prime})$. By construction, $|\sigma|=n$ and $|\mathbf{Zeros}(\sigma)|=z$. Notice that inserting the value 1 in $\sigma$ at any position in $[p_{s-1}+2,n]$ would create an occurrence of 102, since $\sigma_{p_{s-1}+1}=2$ and $\sigma_{n}>2$. It follows that the active sites of $\sigma$ are $\{p_{i}+1\;:\;i\in[1,s-1]\}\sqcup\{n+1\}$, so $|\mathbf{Sites}(\sigma)|=s$. ∎
@@
We know from the definition of $\mathfrak{B}$ that $|\mathfrak{A}_{n}|=\mathfrak{b}_{n}+1$ for all $n\geqslant 0$, so it follows from Corollary 7 that
@@
The formula of Lemma 9 is equivalent to the functional equation
@@
We proceed with the kernel method. The kernel in (3) is $K_{1}(t,u,v)$. This kernel can be cancelled by substituting $v$ for a formal power series $V(t,u)$ which is algebraic with minimal polynomial

### v2

@@ 4 Inversion sequences avoiding 010 and 102
We begin with two simple but useful properties satisfied by inversion sequences avoiding the pattern 010.
@@
4.1 Modifying the construction
As we showed in Proposition 2, the generating tree growing on the left is not always well-defined for inversion sequences avoiding a pattern with several zeros. Specifically, the children of sequences of $\mathcal{I}_{n}(010,102)$ in this tree are precisely the sequences of $\mathcal{I}_{n+1}(010,102)$ in which the value 1 does not appear to the right of a greater value (otherwise their parent sequence would contain a 0 to the right of a greater value, and therefore contain the pattern 010). To complete this construction, we shall introduce some new steps so that the value 1 may appear at any position (while avoiding the patterns 010 and 102).
A possible solution could be to modify the way that entries of value 1 are placed, allowing them to be a newly inserted element in the sequence rather than a replacement of a zero entry. However, we remark that there are some unwanted restrictions on inversion sequences that can be generated by inserting a new 1 in this manner. Indeed, if a sequence $\tau$ were obtained by inserting a new entry of value 1 at position $i$ in an inversion sequence $\sigma\in\mathcal{I}_{n}$, then for all $j\in[i+1,n+1]$, we would have $\tau_{j}=\sigma_{j-1}<j-1$. In particular, a sequence $\tau$ which contains a saturated entry $\tau_{j}=j-1$ for some $j>i$ cannot be generated this way. Nevertheless, for the class $\mathcal{I}(010,102)$, we specifically want to generate a sequence $\tau$ which avoids 102, and in which the value 1 inserted at position $i$ is to the right of a value $k>1$. In particular, for all $j>i$, $\tau$ must satisfy $\tau_{j}\leqslant k<i-1<j-1$ to avoid 102, so the aforementioned problem disappears. Do note that inserting an entry of value 1 to the left of every greater value would still cause this problem, so we cannot place every single entry of value 1 by insertion either. This leads us to consider a hybrid approach in which entries of value 1 are only placed by insertion if they are to the right of a greater value, and still placed by replacing a zero otherwise. This describes a new generating tree for $\mathcal{I}(010,102)$, in which the parent of each nonempty sequence is obtained by
@@ 4.1 Modifying the construction
  • removing the first zero and subtracting 1 from all nonzero entries (as in Section 2), otherwise.
@@
We introduce a second small change in this construction. It is difficult to explain precisely the reason for this choice, but it allows for a simpler presentation of the computation of the generating function of $\mathcal{I}(010,102)$ in Theorem 11, by reducing the number of cases to consider. If a sequence contains an entry of value 1 with no greater value to its left (i.e. the sequence begins with a prefix of zeros, followed by a 1), then its parent is now obtained by subtracting 1 from the leftmost entry of value 1 only [^4] When exploring the ancestors of a sequence of this form, its parent in our original definition will eventually appear, after all entries of value 1 have been turned into zeros. We are simply placing some intermediate sequences on that path.. Otherwise, its parent is determined as described just above. The corresponding tree is represented in Figure 2. It should be noted that this tree does not quite conform to the usual definition of a generating tree, as some nodes are on the same level as their parent.
@@
The article [FPPR03] introduces the notion of “jumping” generating trees, in which each edge has a (bounded) positive integer length. If a node at level $n$ has a child through an edge of length $k$, then this child is at level $n+k$ (in a classical generating tree, all edges have length $1$). The tree in Figure 2 could be seen as a jumping generating tree having edges of length 0 or 1.
We claim that the following succession rule describes a generating tree isomorphic to the one described above (and represented in Figure 2); the arrow $\overset{k}{\leadsto}$ marks an edge of length $k$.
$\Omega_{\{010,102\}}=\left\{\begin{array}[]{rclll}(\mathtt{C},0)\\ (\mathtt{C},z)&\overset{1}{\leadsto}&(\mathtt{C},z+1)\\ &\overset{0}{\leadsto}&(\mathtt{B},z-1,2)&\text{if}&z\geqslant 2\\ (\mathtt{B},z,s)&\overset{1}{\leadsto}&(\mathtt{P},z+1,s)\\ &\overset{0}{\leadsto}&(\mathtt{B},z-1,s+1)&\text{if}&z\geqslant 2\\ (\mathtt{P},z,s)&\overset{1}{\leadsto}&(\mathtt{P},z+1,s)\\ &\overset{1}{\leadsto}&(\mathtt{R},z,s+1)\\ &\overset{1}{\leadsto}&(\mathtt{M},z,i)&\text{for}&i\in[3,s]\\ &\overset{0}{\leadsto}&(\mathtt{L},z-1)\\ (\mathtt{L},z)&\overset{1}{\leadsto}&(\mathtt{P},z+1,2)\\ &\overset{0}{\leadsto}&(\mathtt{L},z-1)&\text{if}&z\geqslant 2\\ (\mathtt{R},z,s)&\overset{1}{\leadsto}&(\mathtt{P},z+1,s)\\ &\overset{1}{\leadsto}&(\mathtt{R},z,s+1)\\ (\mathtt{M},z,s)&\overset{1}{\leadsto}&(\mathtt{P},z+1,s)\\ &\overset{1}{\leadsto}&(\mathtt{M},z,s).\end{array}\right$.
The generating function of $\mathcal{I}(010,102)$ can be computed from the succession rule $\Omega_{\{010,102\}}$, with the kernel method. However, this result is easier to achieve by studying only the subset of sequences of $\mathcal{I}(010,102)$ which do not contain the value 1, and this is how we will proceed in Section 4.2. Due to this restriction, the construction used in Section 4.2 “jumps” over sequences containing the value 1, so they are grayed out in Figure 2.
For the proof of Theorem 11 to come in Section 4.2, it is not necessary to prove our claim that the rule $\Omega_{\{010,102\}}$ describes a generating tree isomorphic to that of Figure 2. Nevertheless, to help convince the reader of the validity of our claim, we now briefly explain the correspondence between inversion sequences and the labels appearing in this rule. In the following descriptions, the letters $\mu$ and $\nu$ denote nonempty sequences of values greater than or equal to 2.
  • The letter $\mathtt{C}$ marks constant sequences of the form $(0)^{a}$ with $a\geqslant 0$.
  • The letter $\mathtt{B}$ marks binary sequences of the form $(0)^{a}\cdot(1)^{b}$ with $a,b>0$.
  • The letter $\mathtt{P}$ marks pushed sequences of the form $(0,\dots,0)\cdot\mu$. These sequences are named after the $\mathbf{push}$ function of Section 4.2, and they correspond to the set $\mathfrak{B}$ in Section 4.2.
  • The letter $\mathtt{L}$ marks sequences whose entries of value 1 are on the left, of the form $(0,\dots,0,1\dots,1)\cdot\mu$.
  • The letter $\mathtt{R}$ marks sequences whose entries of value 1 are on the right, of the form $(0,\dots,0)\cdot\mu\cdot(1,\dots,1)$.
  • The letter $\mathtt{M}$ marks sequences whose entries of value 1 are in the middle, of the form $(0,\dots,0)\cdot\mu\cdot(1,\dots,1)\cdot\nu$.
  • The parameter $z$ counts the zeros of an inversion sequence.
  • The parameter $s$ counts the active sites of an inversion sequence, for some definition of active sites that is similar, albeit slightly different, from that of Section 4.2.
@@
4.2 Enumerating $\mathcal{I}(010,102)$
For all $n\geqslant 0$, let $\mathfrak{A}_{n}=\{\sigma\in\mathcal{I}_{n}(010,102)\;:\;1\notin\mathbf{Vals}(\sigma)\}$ be the set of inversion sequences of size $n$ which avoid 010 and 102, and do not contain the value 1. Let $\mathfrak{A}=\coprod_{n\geqslant 0}\mathfrak{A}_{n}$. For all $\sigma\in\mathcal{I}$, let $\mathbf{push}(\sigma)=0\cdot(\sigma_{i}+\chi(\sigma_{i}>0))_{i\in[0,n]}$ be the sequence obtained by inserting a zero at the beginning of $\sigma$, and adding 1 to the value of all nonzero entries.
Proposition 7.
@@ 4.2 Enumerating $\mathcal{I}(010,102)$
[…]
@@
Let $n\geqslant 0$. The function $\mathbf{push}$ is injective, so it describes a bijection between $\mathcal{I}_{n}(010,102)$ and $\mathbf{push}(\mathcal{I}_{n}(010,102))$. By construction, the image under $\mathbf{push}$ of a sequence of size $n$ is a sequence of size $n+1$, so it follows from Proposition 7 that $\mathbf{push}(\mathcal{I}_{n}(010,102))=\mathfrak{A}_{n+1}$. ∎
@@
If $\sigma$ already contains the value 1, then a new entry of value 1 can only be inserted next to the current entries of value 1, by Proposition 6, and we have little interest in the active sites of $\sigma$ in that case. We next characterize the active sites of $\sigma$ when $\sigma$ does not contain the value 1. Figure 3 represents the position of such an active site, as described in Proposition 9 below.
@@
The value 1 cannot be inserted at positions $[2,z]$, since it would create a subsequence $(0,1,0)$. Inserting the value 1 at any other position (i.e. in $[z+1,n+1])$ cannot create an occurrence of 010, since there are no zeros to the right and no other entry of value $1$ in $\sigma$. It follows that $\mathbf{Sites}(\sigma)$ is the subset of $[z+1,n+1]$ of positions where inserting the value 1 does not create an occurrence of 102. Note that all values of $(\sigma_{i})_{i\in[z+1,n]}$ are greater than 1, so the new entry of value 1 could only take the role of the minimal value in an occurrence of 102. In other words, $i\in[z+1,n+1]$ is an active site of $\sigma$ if and only if no value of $(\sigma_{j})_{j\in[z+1,i-1]}$ is less than a value of $(\sigma_{j})_{j\in[i,n]}$. ∎
[…]
@@
Let $\sigma^{\prime}=\mathbf{push}^{-1}(\sigma)$ be the sequence obtained by removing the first zero from $\sigma$, and subtracting 1 from every nonzero entry. In particular, $\sigma^{\prime}\in\mathcal{I}_{n-1}(010,102)$.
  • 1
@@
    • 2.1
      If $\sigma^{\prime}_{z}=1$, then $\sigma^{\prime}$ is of the form $(0)^{z-1}\cdot(1,\dots,1)\cdot\omega$, where $\omega$ is a sequence of values strictly greater than 1, by Proposition 6.
      • 2.1.1
@@
        If $\omega$ is empty, then $\sigma^{\prime}=(0)^{z-1}\cdot(1)^{n-z}$, so $\sigma=(0)^{z}\cdot(2)^{n-z}$. In that case, $\mathbf{Sites}(\sigma)=[z+1,n+1]$, so $z+s=n+1$.
      • 2.1.2
@@
        If $\omega$ is not empty, we have $\sigma_{z+1}=2$ and $\sigma_{n}>2$. This implies that inserting the value 1 in $\sigma$ at any position in $[z+2,n]$ would create an occurrence of 102. It follows that $\mathbf{Sites}(\sigma)=\{z+1,n+1\}$, so $s=2$.
@@
        Since the factor of ones of $\sigma^{\prime}$ is consecutive to its factor of zeros, replacing all the ones by zeros yields a sequence $\tau=(0,\dots,0)\cdot\omega$, which still avoids 010 and 102. Notice that $\tau$ does not contain the value 1, but contains a value greater than 1, so $\tau\in\mathfrak{B}_{n-1}$. By construction, the number $z^{\prime}$ of zeros of $\tau$ is at least $z$.
    • 2.2
      If $\sigma^{\prime}_{z}>1$ and $\sigma^{\prime}_{n-1}=1$, let $k$ be the number of entries of value $1$ in $\sigma^{\prime}$, and let $\tau=(\sigma^{\prime}_{i})_{i\in[1,n-k-1]}$ be the sequence obtained by removing all entries of value 1 from $\sigma^{\prime}$. In other words, $\sigma^{\prime}=\tau\cdot(1)^{k}$, and $\tau=(0)^{z-1}\cdot\omega$, where $\omega$ is a nonempty sequence of values strictly greater than 1. Notice that $\sigma$ ends with $k$ entries of value $2$, and that inserting the value 1 after any of these entries would not create an occurrence of 102. These $k$ entries (or, equivalently, the rightmost $k$ active sites of $\sigma$) are no longer present $\tau$. This means that $\tau$ has sks-k actives sites, and so $\tau\in\mathfrak{B}_{n-k-1,z-1,s-k}$.
    • 2.3
      If $\sigma^{\prime}_{z}>1$ and $\sigma^{\prime}_{n-1}>1$, then $\sigma^{\prime}$ is of the form $(0)^{z-1}\cdot\mu\cdot(1,\dots,1)\cdot\nu$, where $\mu$ and $\nu$ are two nonempty sequences of values strictly greater than 1, which satisfy $\mathbf{min}(\mu)\geqslant\mathbf{max}(\nu)$ since $\sigma^{\prime}$ avoids 102. Let $p=z+1+|\mu|\in[z+2,n-1]$ be the position of the leftmost entry of value 2 in $\sigma$. By Proposition 9, we have $\{z+1,p,n+1\}\subseteq\mathbf{Sites}(\sigma)$, so $s\geqslant 3$.
[…]
@@
To complete the proof, it remains to show that each of the five cases above induces a bijection between the set of sequences $\sigma$ and the set of sequences $\sigma^{\prime}$ or $\tau$. We now describe the inverse bijection in each case.
  • 1
    Let $n\geqslant 3,z,s\geqslant 2$, $\sigma^{\prime}\in\mathfrak{B}_{n-1,z-1,s}$, and $\sigma=\mathbf{push}(\sigma^{\prime})$. Then $\sigma\in\mathfrak{B}_{n,z,s}$ and does not contain the value 1.
  • 2.1.1
    Let $n\geqslant 3,z\geqslant 2$, $\sigma^{\prime}=(0)^{z-1}\cdot(1)^{n-z}$, and $\sigma=\mathbf{push}(\sigma^{\prime})$. Then $\sigma=(0)^{z}\cdot(2)^{n-z}$.
  • 2.1.2
@@
    Let $n\geqslant 3,z\geqslant 2,z^{\prime}\geqslant z$, and $\tau\in\mathfrak{B}_{n-1,z^{\prime}}$. Let $\sigma^{\prime}$ be the sequence obtained by replacing all but the first z1z-1 zeros of $\tau$ by ones, and let $\sigma=\mathbf{push}(\sigma^{\prime})$. Note that the transformation from $\tau$ to $\sigma^{\prime}$ cannot create any occurrence of 010 or 102, since there are no zeros to the right of the ones in $\sigma^{\prime}$. It follows from Proposition 7 that $\sigma$ still avoids 010 and 102. By construction, $\sigma_{z+1}=\sigma^{\prime}_{z}+1=2$, and $\sigma_{n}=\tau_{n-1}+1>2$, hence inserting the value 1 in $\sigma$ at any position in $[z+2,n]$ would create an occurrence of 102. It follows that $z+1$ and $n+1$ are the only two active sites of $\sigma$, so $\sigma\in\mathfrak{B}_{n,z,2}$.
  • 2.2
    Let $n\geqslant 3,z,s\geqslant 2,k\geqslant 1$, and $\tau\in\mathfrak{B}_{n-k-1,z-1,s-k}$. Let $\sigma^{\prime}=\tau\cdot(1)^{k}$, and let $\sigma=\mathbf{push}(\sigma^{\prime})$. By Proposition 9, inserting the value 1 at the end of a sequence cannot create an occurrence of 010 or 102, so $\sigma$ still avoids both patterns. It can be seen that $\sigma$ ends with $k$ entries of value 2, and inserting the value 1 to the right of such an entry cannot create an occurrence of 102, hence all positions in [nk+2,n+1][n-k+2,n+1] are active sites of $\sigma$. Adding to that the sks-k active sites from $\tau$, $\sigma$ has a total of $s$ actives sites, so $\sigma\in\mathfrak{B}_{n,z,s}$.
  • 2.3
    Let $n,s\geqslant 3$, $z\geqslant 2$, $n^{\prime}\leqslant n-2$, $s^{\prime}\geqslant s$, and $\tau\in\mathfrak{B}_{n^{\prime},z-1,s^{\prime}}$. Let $p_{1}<p_{2}<\dots<p_{s^{\prime}}$ be the active sites of $\tau$. Let $\sigma^{\prime}$ be the sequence obtained by inserting $n-n^{\prime}-1$ entries of value 1 at position $p_{s-1}$ in $\tau$, and let $\sigma=\mathbf{push}(\sigma^{\prime})$. Inserting the value 1 in $\sigma$ at any position in $p_{s-1}+2,n]$ would create an occurrence of 102, since $\sigma_{p_{s-1}+1}=2$ and $\sigma_{n}>2$. It follows that the active sites of $\sigma$ are $\{p_{i}+1\;:\;i\in[1,s-1]\}\sqcup\{n+1\}$, so $\sigma\in\mathfrak{B}_{n,z,s}$. ∎
@@
By the definition of $\mathfrak{B}$, we have $|\mathfrak{A}_{n}|=\mathfrak{b}_{n}+1$ for all $n\geqslant 0$, so it follows from Corollary 8 that
@@
By Lemma 10, we have
$B(t,u,v)=\sum_{\begin{subarray}{c}n\geqslant 3\\ z,s\geqslant 2\end{subarray}}$ $\left[\mathfrak{b}_{n-1,z-1,s}+\chi(z+s=n+1)+\chi(s=2)\cdot\left(\sum_{z^{\prime}\geqslant z}\mathfrak{b}_{n-1,z^{\prime}}\right)\right$.
$+\left.\left(\sum_{k\geqslant 1}\mathfrak{b}_{n-k-1,z-1,s-k}\right)+\chi(s\geqslant 3)\cdot\left(\sum_{n^{\prime}\leqslant n-2}\sum_{s^{\prime}\geqslant s}\mathfrak{b}_{n^{\prime},z-1,s^{\prime}}\right)\right]t^{n}u^{z}v^{s}.$
We now simplify each of the five terms in this expression to obtain a functional equation characterizing $B(t,u,v)$.
$\sum_{\begin{subarray}{c}n\geqslant 3\\ z,s\geqslant 2\end{subarray}}\mathfrak{b}_{n-1,z-1,s}t^{n}u^{z}v^{s}=tuB(t,u,v)$
$\sum_{\begin{subarray}{c}n\geqslant 3\\ z,s\geqslant 2\end{subarray}}\chi(z+s=n+1)t^{n}u^{z}v^{s}=\sum_{z,s\geqslant 2}t^{-1}(tu)^{z}(tv)^{s}=\frac{t^{3}u^{2}v^{2}}{(1-tu)(1-tv)}$
$\sum_{\begin{subarray}{c}n\geqslant 3\\ z,s\geqslant 2\end{subarray}}\chi(s=2)\cdot\left(\sum_{z^{\prime}\geqslant z}\mathfrak{b}_{n-1,z^{\prime}}\right)t^{n}u^{z}v^{s}$ $=\sum_{\begin{subarray}{c}n\geqslant 2\\ z\geqslant 2\end{subarray}}\sum_{i\geqslant 0}\mathfrak{b}_{n,z+i}t^{n+1}u^{z}v^{2}$
$=\sum_{\begin{subarray}{c}n\geqslant 2\\ z\geqslant 2\end{subarray}}\sum_{i\in[0,z-2]}\mathfrak{b}_{n,z}t^{n+1}u^{z-i}v^{2}$
$=tv^{2}\sum_{\begin{subarray}{c}n\geqslant 2\\ z\geqslant 2\end{subarray}}\mathfrak{b}_{n,z}t^{n}\cdot\frac{u^{2}-u^{z+1}}{1-u}$
$=tv^{2}\left(\frac{u^{2}B(t,1,1)-uB(t,u,1)}{1-u}\right)$
$\sum_{\begin{subarray}{c}n\geqslant 3\\ z,s\geqslant 2\end{subarray}}\left(\sum_{k\geqslant 1}\mathfrak{b}_{n-k-1,z-1,s-k}\right)t^{n}u^{z}v^{s}$ $=\sum_{\begin{subarray}{c}z\geqslant 1\\ n,s\geqslant 2\end{subarray}}\sum_{k\geqslant 1}\mathfrak{b}_{n,z,s}t^{n+k+1}u^{z+1}v^{s+k}$
$=\sum_{\begin{subarray}{c}z\geqslant 1\\ n,s\geqslant 2\end{subarray}}\mathfrak{b}_{n,z,s}t^{n+1}u^{z+1}v^{s}\sum_{k\geqslant 1}(tv)^{k}$
$=\sum_{\begin{subarray}{c}z\geqslant 1\\ n,s\geqslant 2\end{subarray}}\mathfrak{b}_{n,z,s}t^{n+1}u^{z+1}v^{s}\cdot\frac{tv}{1-tv}$
$=t^{2}uv\left(\frac{B(t,u,v)}{1-tv}\right)$
$\sum_{\begin{subarray}{c}n\geqslant 3\\ z,s\geqslant 2\end{subarray}}\chi(s\geqslant 3)\cdot\left(\sum_{n^{\prime}\leqslant n-2}\sum_{s^{\prime}\geqslant s}\mathfrak{b}_{n^{\prime},z-1,s^{\prime}}\right)t^{n}u^{z}v^{s}$ $=\sum_{\begin{subarray}{c}n,s\geqslant 3\\ z\geqslant 1\end{subarray}}\sum_{\begin{subarray}{c}i\geqslant 2\\ j\geqslant 0\end{subarray}}\mathfrak{b}_{n-i,z,s+j}t^{n}u^{z+1}v^{s}$
$=\sum_{\begin{subarray}{c}n,s\geqslant 3\\ z\geqslant 1\end{subarray}}\mathfrak{b}_{n,z,s}\sum_{\begin{subarray}{c}i\geqslant 2\\ j\in[0,s-3]\end{subarray}}t^{n+i}u^{z+1}v^{s-j}$
$=\sum_{\begin{subarray}{c}n,s\geqslant 3\\ z\geqslant 1\end{subarray}}\mathfrak{b}_{n,z,s}t^{n}u^{z+1}\cdot\frac{t^{2}}{1-t}\cdot\frac{v^{3}-v^{s+1}}{1-v}$
$=t^{2}u\left(\frac{v^{3}B(t,u,1)-vB(t,u,v)}{(1-t)(1-v)}\right)$
In summary, the formula of Lemma 10 is equivalent to the functional equation
@@
We apply the kernel method. The kernel in (3) is $K_{1}(t,u,v)$. This kernel can be cancelled by substituting $v$ for a formal power series $V(t,u)$ which is algebraic with minimal polynomial